OpenStudy (anonymous):

Solve the system of equations x + 3y = 32 -3x + 2y = 3 Please help me solve this

OpenStudy (anonymous):

x=-3y+32 -3(-3y+32)+2y=3

OpenStudy (anonymous):

Good answer can u explain and show work?

OpenStudy (anonymous):

x + 3y = 32 multiply all of this by 3 then add to the other equation] 3x + 9y = 96 -3x + 2y = 3 [adding equations gives] 11y = 99 [dividing the equation by 11] y = 9 If y = 9 substitute into frst equation to solve for x x + 3(9) = 32 x + 27 = 32 [subtract 27 from both sides] x - 5 Answer x = 5 , y = 9

OpenStudy (anonymous):

or as imranmeah did, first one for x and then substitute into second one \[x=-3y+32 \] \[-3(-3y+32)+2y=3\]

OpenStudy (anonymous):

thank u everyone-i have 2 more i need help with Best Rentals charges a daily fee plus a mileage fee for renting its cars. Barney was charged $123.00 for 3 days and 300 miles, while Mary was charged $216.00 for 5 days and 600 miles. What does Best Rental charge per day, for mileage?

OpenStudy (anonymous):

ok so i still need to solve ?